OpenStudy (anonymous):

Find the sum of the geometric sequence 1, 1/4, 1/16, 1/64, 1/256 show all work please

OpenStudy (kainui):

S is our sum, x is 1/4 in this case, and we're adding them up to some power of 1/4. It looks like n=4 because (1/4)^4=1/256. \[\Large S= x^0+x^1+x^2+...+x^n \\ \Large S-x^0= x^1+x^2+...+x^n \\ \Large \frac{S-x^0}{x^1}= x^0+x^1+...+x^{n-1} \\ \Large \frac{S-x^0}{x^1}+x^n= x^0+x^1+...+x^{n-1}+ x^n \\ \Large \frac{S-x^0}{x^1}+x^n=S\] Now we can solve for S. \[\Large \frac{S-1}{x}+x^n=S\]

OpenStudy (anonymous):

Thank You this helps a lot!

OpenStudy (kropot72):

You might find this formula useful: \[S _{n}=\frac{a(1-r ^{n})}{1-r}\] which is the sum of n terms of a geometric series with first term a and common ratio r.